### v1

# Classic GNNs are Strong Baselines: Reassessing GNNs for Node Classification

Abstract
Graph Transformers (GTs) have recently emerged as popular alternatives to traditional message-passing Graph Neural Networks (GNNs), due to their theoretically superior expressiveness and impressive performance reported on standard node classification benchmarks, often significantly outperforming GNNs. In this paper, we conduct a thorough empirical analysis to reevaluate the performance of three classic GNN models (GCN, GAT, and GraphSAGE) against GTs. Our findings suggest that the previously reported superiority of GTs may have been overstated due to suboptimal hyperparameter configurations in GNNs. Remarkably, with slight hyperparameter tuning, these classic GNN models achieve state-of-the-art performance, matching or even exceeding that of recent GTs across 17 out of the 18 diverse datasets examined. Additionally, we conduct detailed ablation studies to investigate the influence of various GNN configurations—such as normalization, dropout, residual connections, network depth, and jumping knowledge mode—on node classification performance. Our study aims to promote a higher standard of empirical rigor in the field of graph machine learning, encouraging more accurate comparisons and evaluations of model capabilities. Our implementation is available at https://github.com/LUOyk1999/tunedGNN.
1 Introduction
Node classification is a fundamental task in graph machine learning, with high-impact applications across many fields such as social network analysis, bioinformatics, and recommendation systems. Graph Neural Networks (GNNs)  [20, 28, 65, 72, 49, 33, 8, 76, 18, 9, 51, 4, 73, 56, 10, 52, 53, 66, 77, 37] have emerged as a powerful class of models for tackling the node classification task. GNNs operate by iteratively aggregating information from a node’s neighbors, a process known as message passing [19], leveraging both the graph structure and node features to learn useful node representations for classification. While GNNs have achieved notable success, studies have identified several limitations, including over-smoothing [35], over-squashing [1], lack of sensitivity to heterophily [81], and challenges in capturing long-range dependencies [11].
Recently, Graph Transformers (GTs) [50, 48, 23] have gained prominence as popular alternatives to GNNs. Unlike GNNs, which primarily aggregate local neighborhood information, the Transformer architecture [64] can capture interactions between any pair of nodes via a self-attention layer. GTs have achieved significant success in graph-level tasks, e.g., graph classification involving small-scale graphs like molecular graphs [13, 74, 30, 42, 55, 6]. This success has inspired efforts [12, 17, 71, 70, 69, 79, 82, 15, 60, 7, 29, 40, 38] to utilize GTs to tackle node classification tasks, especially on large-scale graphs, addressing the aforementioned limitations of GNNs. While recent advancements in state-of-the-art GTs [12, 71] have shown promising results, it’s observed that many of these models, whether explicitly or implicitly, still rely on GNNs for learning local node representations, integrating them alongside the global attention mechanisms for a more comprehensive representation.
This prompts us to reconsider: Could the potential of message-passing GNNs for node classification have been previously underestimated? While prior research has addressed this issue to some extent [24, 14, 68, 44, 54], these studies have limitations in terms of scope and comprehensiveness, including a restricted number and diversity of datasets, as well as an incomplete examination of hyperparameters. In this study, we comprehensively reassess the performance of GNNs for node classification, utilizing three classic GNN models—GCN [28], GAT [64], and GraphSAGE [20]—across 18 real-world benchmark datasets that include homophilous, heterophilous, and large-scale graphs. We examine the influence of key hyperparameters on GNN training, including normalization [2, 26], dropout [63], residual connections [21], network depth, and the utilization of the jumping knowledge mode [73]. We summarize the key findings in our empirical study as follows:
  • With proper hyperparameter tuning, classic GNNs can achieve highly competitive performance in node classification across homophilous and heterophilous graphs with up to millions of nodes. Notably, classic GNNs outperform state-of-the-art GTs, achieving the top rank on 17 out of 18 datasets. This indicates that the previously claimed superiority of GTs over GNNs may have been overstated, possibly due to suboptimal hyperparameter configurations in GNN evaluations.
  • Our ablation studies have yielded valuable insights into GNN hyperparameters for node classification. We demonstrate that (1) normalization is essential for large-scale graphs; (2) dropout consistently proves beneficial; (3) residual connections can significantly enhance performance, especially on heterophilous graphs; (4) GNNs on heterophilous graphs tend to perform better with deeper layers; and (5) jumping knowledge connections have minimal impact on large-scale graphs.
2 Classic GNNs for Node Classification
Define a graph as $\mathcal{G}=(\mathcal{V},\mathcal{E},\boldsymbol{X},\boldsymbol{Y})$, where $\mathcal{V}$ denotes the set of nodes, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ represents the set of edges, $\boldsymbol{X}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is the node feature matrix, with $|\mathcal{V}|$ representing the number of nodes and $d$ the dimension of the node features, and $\boldsymbol{Y}\in\mathbb{R}^{|\mathcal{V}|\times C}$ is the one-hot encoded label matrix, with $C$ being the number of classes. Let $\boldsymbol{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ denote the adjacency matrix of $\mathcal{G}$.
Message Passing Graph Neural Networks (GNNs) [19] compute node representations $\boldsymbol{h}_{v}^{l}$ at each layer $l$ as:
$\boldsymbol{h}_{v}^{l}=\text{UPDATE}^{l}\left(\boldsymbol{h}_{v}^{l-1},\text{AGG}^{l}\left(\left\{\boldsymbol{h}_{u}^{l-1}\mid u\in\mathcal{N}\left(v\right)\right\}\right)\right),$ (1)
where $\mathcal{N}(v)$ represents the neighboring nodes adjacent to $v$, $\text{AGG}^{l}$ serves as the message aggregation function, and $\text{UPDATE}^{l}$ is the update function. Initially, each node $v$ begins with a feature vector $\boldsymbol{h}_{v}^{0}=\boldsymbol{x}_{v}\in\mathbb{R}^{d}$. The function $\text{AGG}^{l}$ aggregates information from the neighbors of $v$ to update its representation. The output of the last layer $L$, i.e., $\text{GNN}(v,\boldsymbol{A},\boldsymbol{X})=\boldsymbol{h}_{v}^{L}$, is the representation of $v$ produced by the GNN. In this work, we focus on three classic GNNs: GCN [28], GraphSAGE [20], and GAT [64], which differ in their approach to learning the node representation $\boldsymbol{h}_{v}^{l}$.
Graph Convolutional Networks (GCN) [28], the standard GCN model, is formualated as:
$\boldsymbol{h}_{v}^{l}=\sigma(\sum_{u\in\mathcal{N}(v)\cup\{v\}}\frac{1}{\sqrt{\hat{d}_{u}\hat{d}_{v}}}\boldsymbol{h}_{u}^{l-1}\boldsymbol{W}^{l}),$ (2)
where $\hat{d}_{v}=1+\sum_{u\in\mathcal{N}(v)}1$, $\sum_{u\in\mathcal{N}(v)}1$ denotes the degree of node $v$, $\boldsymbol{W}^{l}$ is the trainable weight matrix in layer $l$, and $\sigma$ is the activation function, e.g., ReLU(·) = $\max(0,\text{·})$.
GraphSAGE [20] learns node representations through a different approach:
$\boldsymbol{h}_{v}^{l}=\sigma(\boldsymbol{h}_{v}^{l-1}\boldsymbol{W}_{1}^{l}+(\text{mean}_{u\in\mathcal{N}(v)}\boldsymbol{h}_{u}^{l-1})\boldsymbol{W}_{2}^{l}),$ (3)
where $\boldsymbol{W}_{1}^{l}$ and $\boldsymbol{W}_{2}^{l}$ are trainable weight matrices, and $\text{mean}_{u\in\mathcal{N}(v)}\boldsymbol{h}_{u}^{l-1}$ computes the average embedding of the neighboring nodes of $v$.
Graph Attention Networks (GAT) [64] employ masked self-attention to assign weights to different neighboring nodes. For an edge $(v,u)\in\mathcal{E}$, the propagation rule of GAT is defined as:
$\alpha_{vu}^{l}=\frac{\exp\left(\text{LeakyReLU}\left(\mathbf{a}^{\top}_{l}\left[\boldsymbol{W}^{l}\boldsymbol{h}_{v}^{l-1}\,\|\,\boldsymbol{W}^{l}\boldsymbol{h}_{u}^{l-1}\right]\right)\right)}{\sum_{r\in\mathcal{N}(v)}\exp\left(\text{LeakyReLU}\left(\mathbf{a}^{\top}_{l}\left[\boldsymbol{W}^{l}\boldsymbol{h}_{v}^{l-1}\,\|\,\boldsymbol{W}^{l}\boldsymbol{h}_{r}^{l-1}\right]\right)\right)},$
$\boldsymbol{h}_{v}^{l}=\sigma\left(\sum_{u\in\mathcal{N}(v)}\alpha_{vu}^{l}\boldsymbol{h}_{u}^{l-1}\boldsymbol{W}^{l}\right),$ (4)
[…]
Node Classification aims to predict the labels of the unlabeled nodes. Typically, for any node $v$, the node representation generated by the last GNN layer is passed through a prediction head $g(\text{·})$, to obtain the predicted label $\hat{\boldsymbol{y}}_{v}=g(\text{GNN}(v,\boldsymbol{A},\boldsymbol{X}))$. The training objective is to minimize the total loss $L(\boldsymbol{\theta})=\sum_{v\in\mathcal{V}_{\text{train}}}\ell(\hat{\boldsymbol{y}}_{v},\boldsymbol{y}_{v})$ w.r.t. all nodes in the training set $\mathcal{V}_{\text{train}}$, where $\boldsymbol{y}_{v}$ indicates the ground-truth label of $v$ and $\boldsymbol{\theta}$ indicates the trainable GNN parameters.
Homophilous and Heterophilous Graphs. Node classification can be performed on both homophilous and heterophilous graphs. Homophilous graphs are characterized by edges that tend to connect nodes of the same class, while in heterophilous graphs, connected nodes may belong to different classes [54]. GNN models implicitly assume homophily in graphs [45], and it is commonly believed that due to this homophily assumption, GNNs cannot generalize well to heterophilous graphs [81, 9]. However, recent works [43, 39, 54, 41] have empirically shown that standard GCNs also work well on heterophilous graphs. In this study, we provide a comprehensive evaluation of classic GNNs for node classification on both homophilous and heterophilous graphs.
3 Key Hyperparameters for Training GNNs
In this section, we present an overview of the key hyperparameters for training GNNs, including normalization, dropout, residual connections, network depth, and “jumping knowledge” mode. The first four hyperparameters are widely utilized across different types of neural networks to improve model performance, while the last one is specific to GNNs.
Normalization. Specifically, Layer Normalization (LN) [2] or Batch Normalization (BN) [26] can be used in every layer before the activation function $\sigma(\cdot)$. Taking GCN as an example:
$\boldsymbol{h}_{v}^{l}=\sigma(\text{Norm}(\sum_{u\in\mathcal{N}(v)\cup\{v\}}\frac{1}{\sqrt{\hat{d}_{u}\hat{d}_{v}}}\boldsymbol{h}_{u}^{l-1}\boldsymbol{W}^{l})).$ (5)
The normalization techniques are essential for stabilizing the training process by reducing the covariate shift, which occurs when the distribution of each layer’s node embeddings changes during training. Normalizing the node embeddings helps to maintain a more consistent distribution, allowing the use of higher learning rates and leading to faster convergence [5].
Dropout [63], a technique widely used in convolutional neural networks (CNNs) to address overfitting by reducing co-adaptation among hidden neurons [22, 75], has also been found to be effective in addressing similar issues in GNNs [64, 61], where the co-adaptation effects propagate and accumulate through message passing among different nodes. Typically, dropout is applied to the feature embeddings after the activation function:
$\boldsymbol{h}_{v}^{l}=\text{Dropout}(\sigma(\text{Norm}(\sum_{u\in\mathcal{N}(v)\cup\{v\}}\frac{1}{\sqrt{\hat{d}_{u}\hat{d}_{v}}}\boldsymbol{h}_{u}^{l-1}\boldsymbol{W}^{l}))).$ (6)
Residual Connections [21] significantly enhance CNN performance by connecting layer inputs directly to outputs, thereby alleviating the vanishing gradient issue. They were first adopted by the seminal GCN paper [28] and subsequently incorporated into DeepGCNs [33] to boost performance. Formally, linear residual connections can be integrated into GNNs as follows:
$\boldsymbol{h}_{v}^{l}=\text{Dropout}(\sigma(\text{Norm}(\boldsymbol{h}_{u}^{l-1}\boldsymbol{W_{r}}^{l}+\sum_{u\in\mathcal{N}(v)\cup\{v\}}\frac{1}{\sqrt{\hat{d}_{u}\hat{d}_{v}}}\boldsymbol{h}_{u}^{l-1}\boldsymbol{W}^{l}))),$ (7)
where $\boldsymbol{W}_{r}^{l}$ is a trainable weight matrix. This configuration mitigates gradient instabilities and enhances GNN expressiveness [72], addressing the over-smoothing [35] and oversquashing [1] issues since the linear component ($\boldsymbol{h}_{u}^{l-1}\boldsymbol{W_{r}}^{l}$) helps to preserve distinguishable node representations [68].
Network Depth. Deeper network architectures, such as deep CNNs [21, 25], are capable of extracting more complex, high-level features from data, potentially leading to better performance on various prediction tasks. However, GNNs face unique challenges with depth, such as over-smoothing [35], where node representations become indistinguishable with increased network depth. Consequently, in practice, most GNNs adopt a shallow architecture, typically consisting of 2 to 5 layers. While previous research, such as DeepGCN [33] and DeeperGCN [34], advocates the use of deep GNNs with up to 56 and 112 layers, our findings indicate that comparable performance can be achieved with significantly shallower GNN architectures, typically ranging from 2 to 10 layers.
Jumping Knowledge (JK) Mode [73] aggregates representations from different GNN layers, effectively capturing information from varying neighborhood ranges within the graph. For any node $v$, the summation version of JK mode produces the representation of $v$ by:
$\text{GNN}_{\text{JK}}(v,\boldsymbol{A},\boldsymbol{X})=\boldsymbol{h}_{v}^{1}+\boldsymbol{h}_{v}^{2}+\ldots+\boldsymbol{h}_{v}^{L},$ (8)
where $L$ is the number of GNN layers.
4 Experimental Setup for Node Classification
Table 1: Overview of the datasets used for node classification.
Dataset Type # nodes # edges # Features Classes Metric
Cora Homophily 2,708 5,278 2,708 7 Accuracy
CiteSeer Homophily 3,327 4,522 3,703 6 Accuracy
PubMed Homophily 19,717 44,324 500 3 Accuracy
Computer Homophily 13,752 245,861 767 10 Accuracy
Photo Homophily 7,650 119,081 745 8 Accuracy
CS Homophily 18,333 81,894 6,805 15 Accuracy
Physics Homophily 34,493 247,962 8,415 5 Accuracy
WikiCS Homophily 11,701 216,123 300 10 Accuracy
Squirrel Heterophily 5,201 216,933 2,089 5 Accuracy
Chameleon Heterophily 2,277 36,101 2,325 5 Accuracy
Roman-Empire Heterophily 22,662 32,927 300 18 Accuracy
Amazon-Ratings Heterophily 24,492 93,050 300 5 Accuracy
Minesweeper Heterophily 10,000 39,402 7 2 ROC-AUC
Questions Heterophily 48,921 153,540 301 2 ROC-AUC
ogbn-proteins Homophily (Large graphs) 132,534 39,561,252 8 2 ROC-AUC
ogbn-arxiv Homophily (Large graphs) 169,343 1,166,243 128 40 Accuracy
ogbn-products Homophily (Large graphs) 2,449,029 61,859,140 100 47 Accuracy
pokec Heterophily (Large graphs) 1,632,803 30,622,564 65 2 Accuracy
Datasets. Table 1 presents a summary of the statistics and characteristics of the datasets.
  • Homophilous Graphs. Cora, CiteSeer, and PubMed are three widely used citation networks [58]. We follow the semi-supervised setting of [28] for data splits and metrics. Additionally, Computer and Photo [59] are co-purchase networks where nodes represent goods and edges indicate that the connected goods are frequently bought together. CS and Physics [59] are co-authorship networks where nodes denote authors and edges represent that the authors have co-authored at least one paper. We adhere to the widely accepted practice of training/validation/test splits of 60%/20%/20% and metric of accuracy [7, 60, 12]. Furthermore, we utilize the WikiCS dataset and use the official splits and metrics provided in [47].
  • Heterophilous Graphs. Squirrel and Chameleon [57] are two well-known page-page networks that focus on specific topics in Wikipedia. According to the heterophilous graphs benchmarking paper [54], the original split of these datasets introduces overlapping nodes between training and testing, leading to the proposal of a new data split that filters out the overlapping nodes. We use its provided split and its metrics for evaluation. Additionally, we utilize four other heterophilous datasets proposed by the same source [54]: Roman-Empire, where nodes correspond to words in the Roman Empire Wikipedia article and edges connect sequential or syntactically linked words; Amazon-Ratings, where nodes represent products and edges connect frequently co-purchased items; Minesweeper, a synthetic dataset where nodes are cells in a $100\times 100$ grid and edges connect neighboring cells; and Questions, where nodes represent users from the Yandex Q question-answering website and edges connect users who interacted through answers. All splits and evaluation metrics are consistent with those proposed in the source.
  • Large-scale Graphs. We consider a collection of large graphs released recently by the Open Graph Benchmark (OGB) [24]: ogbn-arxiv, ogbn-proteins, and ogbn-products, with node numbers ranging from 0.16M to 2.4M. We maintain all the OGB standard evaluation settings. Additionally, we analyze performance on the social network pokec [32], which has 1.6M nodes, following the evaluation settings of [12].
Table 2: Node classification results over homophilous graphs (%). indicates our implementation, while other results are taken from [12, 71]. The top $\mathbf{1^{st}}$, $\mathbf{2^{nd}}$ and $\mathbf{3^{rd}}$ results are highlighted.
Cora CiteSeer PubMed Computer Photo CS Physics WikiCS
# nodes 2,708 3,327 19,717 13,752 7,650 18,333 34,493 11,701
# edges 5,278 4,732 44,324 245,861 119,081 81,894 247,962 216,123
[…]
GraphGPS 82.84 ± 1.03 72.73 ± 1.23 79.94 ± 0.26 91.19 ± 0.54 95.06 ± 0.13 93.93 ± 0.12 97.12 ± 0.19 78.66 ± 0.49
NAGphormer 82.12 ± 1.18 71.47 ± 1.30 79.73 ± 0.28 91.22 ± 0.14 95.49 ± 0.11 95.75 ± 0.09 97.34 ± 0.03 77.16 ± 0.72
Exphormer 82.77 ± 1.38 71.63 ± 1.19 79.46 ± 0.35 91.47 ± 0.17 95.35 ± 0.22 94.93 ± 0.01 96.89 ± 0.09 78.54 ± 0.49
GOAT 83.18 ± 1.27 71.99 ± 1.26 79.13 ± 0.38 90.96 ± 0.90 92.96 ± 1.48 94.21 ± 0.38 96.24 ± 0.24 77.00 ± 0.77
NodeFormer 82.20 ± 0.90 72.50 ± 1.10 79.90 ± 1.00 86.98 ± 0.62 93.46 ± 0.35 95.64 ± 0.22 96.45 ± 0.28 74.73 ± 0.94
SGFormer 84.50 ± 0.80 72.60 ± 0.20 80.30 ± 0.60 91.99 ± 0.76 95.10 ± 0.47 94.78 ± 0.20 96.60 ± 0.18 73.46 ± 0.56
Polynormer 83.25 ± 0.93 72.31 ± 0.78 79.24 ± 0.43 93.68 ± 0.21 96.46 ± 0.26 95.53 ± 0.16 97.27 ± 0.08 80.10 ± 0.67
GCN 81.60 ± 0.40 71.60 ± 0.40 78.80 ± 0.60 89.65 ± 0.52 92.70 ± 0.20 92.92 ± 0.12 96.18 ± 0.07 77.47 ± 0.85
GCN 85.08 ± 0.52 3.48$\uparrow$ 72.98 ± 0.84 1.38$\uparrow$ 81.32 ± 0.72 2.52$\uparrow$ 93.80 ± 0.29 4.15$\uparrow$ 96.51 ± 0.20 3.81$\uparrow$ 95.80 ± 0.28 2.88$\uparrow$ 97.43 ± 0.05 1.25$\uparrow$ 80.27 ± 0.71 2.80$\uparrow$
GraphSAGE 82.68 ± 0.47 71.93 ± 0.85 79.41 ± 0.53 91.20 ± 0.29 94.59 ± 0.14 93.91 ± 0.13 96.49 ± 0.06 74.77 ± 0.95
GraphSAGE 84.18 ± 0.81 1.50$\uparrow$ 71.93 ± 0.85 0.00$\uparrow$ 79.41 ± 0.53 0.00$\uparrow$ 93.59 ± 0.22 2.39$\uparrow$ 96.41 ± 0.17 1.82$\uparrow$ 96.12 ± 0.24 2.21$\uparrow$ 97.21 ± 0.05 0.72$\uparrow$ 80.51 ± 0.48 5.74$\uparrow$
GAT 83.00 ± 0.70 72.10 ± 1.10 79.00 ± 0.40 90.78 ± 0.13 93.87 ± 0.11 93.61 ± 0.14 96.17 ± 0.08 76.91 ± 0.82
GAT 84.64 ± 1.27 1.64$\uparrow$ 72.10 ± 1.10 0.00$\uparrow$ 79.70 ± 0.70 0.70$\uparrow$ 93.93 ± 0.16 3.15$\uparrow$ 96.67 ± 0.13 2.80$\uparrow$ 96.08 ± 0.10 2.47$\uparrow$ 97.30 ± 0.06 1.13$\uparrow$ 80.75 ± 0.74 3.84$\uparrow$
Baselines. Our main focus lies on classic GNNs: GCN [28], GraphSAGE [20], GAT [64], the state-of-the-art scalable GTs: SGFormer [71], Polynormer [12], GOAT [29], NodeFormer [70], NAGphormer [7], and powerful GTs: GraphGPS [55] and Exphormer [60]. Furthermore, various other GTs like [17, 15, 38, 78, 31, 3, 6, 74, 13] exist in related surveys [23, 50], empirically shown to be inferior to the GTs we compared against for node classification tasks. For heterophilous graphs, We also consider five models designed for node classification under heterophily following [54]: H2GCN [81], CPGNN [80], GPRGNN [9], FSGNN [46], GloGNN [36]. Note that we adopt the empirically optimal Polynormer variant (Polynormer-r), which demonstrates superior performance over advanced GNNs such as LINKX [37] and OrderedGNN [62]. We report the performance results of baselines primarily from [12, 71, 54], with the remaining obtained from their respective original papers or official leaderboards whenever possible, as those results are obtained by well-tuned models.
Hyperparameter Configurations. We conduct minimal hyperparameter tuning on classic GNNs, consistent with the parameter search range of Polynormer [12]. We utilize the Adam optimizer [27] with a learning rate from $\{0.001,0.005,0.01\}$ and an epoch limit of 2500. And we tune the hidden dimension from $\{64,256,512\}$. As discussed in Section 3, we focus on whether to use normalization (BN or LN), residual connections, jumping knowledge, and dropout rates from $\{0.2,0.3,0.5,0.7\}$, the number of layers from $\{2,3,4,5,6,7,8,9,10\}$. Considering the large number of hyperparameters and datasets, we do not perform an exhaustive search. We report mean scores and standard deviations after 5 runs. GNN denotes our implementation of the GNN model.
Detailed experimental setup and hyperparameters are provided in Appendix A.
5 Empirical Findings
5.1 Performance of Classic GNNs in Node Classification
In this subsection, we provide a detailed analysis of the performance of the three classic GNNs compared to state-of-the-art GTs in node classification tasks. Our experimental results across homophilous (Table 2), heterophilous (Table 3), and large-scale graphs (Table 4) reveal that classic GNNs often outperform or match the performance of advanced GTs across 18 datasets. Notably, among the 18 datasets evaluated, classic GNNs achieve the top rank on 17 of them, showcasing their robust competitiveness. We highlight our main observations below.
Table 3: Node classification results on heterophilous graphs (%). indicates our implementation, while other results are taken from [12, 71, 54]. The top $\mathbf{1^{st}}$, $\mathbf{2^{nd}}$ and $\mathbf{3^{rd}}$ results are highlighted.
Squirrel Chameleon Amazon-Ratings Roman-Empire Minesweeper Questions
# nodes 2223 890 24,492 22,662 10,000 48,921
# edges 46,998 8,854 93,050 32,927 39,402 153,540
[…]
H2GCN 35.10 ± 1.15 26.75 ± 3.64 36.47 ± 0.23 60.11 ± 0.52 89.71 ± 0.31 63.59 ± 1.46
CPGNN 30.04 ± 2.03 33.00 ± 3.15 39.79 ± 0.77 63.96 ± 0.62 52.03 ± 5.46 65.96 ± 1.95
GPRGNN 38.95 ± 1.99 39.93 ± 3.30 44.88 ± 0.34 64.85 ± 0.27 86.24 ± 0.61 55.48 ± 0.91
FSGNN 35.92 ± 1.32 40.61 ± 2.97 52.74 ± 0.83 79.92 ± 0.56 90.08 ± 0.70 78.86 ± 0.92
GloGNN 35.11 ± 1.24 25.90 ± 3.58 36.89 ± 0.14 59.63 ± 0.69 51.08 ± 1.23 65.74 ± 1.19
GraphGPS 39.67 ± 2.84 40.79 ± 4.03 53.10 ± 0.42 82.00 ± 0.61 90.63 ± 0.67 71.73 ± 1.47
NodeFormer 38.52 ± 1.57 34.73 ± 4.14 43.86 ± 0.35 64.49 ± 0.73 86.71 ± 0.88 74.27 ± 1.46
SGFormer 41.80 ± 2.27 44.93 ± 3.91 48.01 ± 0.49 79.10 ± 0.32 90.89 ± 0.58 72.15 ± 1.31
Polynormer 40.87 ± 1.96 41.82 ± 3.45 54.81 ± 0.49 92.55 ± 0.37 97.46 ± 0.36 78.92 ± 0.89
GCN 38.67 ± 1.84 41.31 ± 3.05 48.70 ± 0.63 73.69 ± 0.74 89.75 ± 0.52 76.09 ± 1.27
GCN 44.50 ± 1.92 5.83$\uparrow$ 46.11 ± 3.16 4.80$\uparrow$ 53.57 ± 0.32 4.87$\uparrow$ 91.35 ± 0.37 17.68$\uparrow$ 97.77 ± 0.38 8.02$\uparrow$ 77.40 ± 1.07 1.31$\uparrow$
GraphSAGE 36.09 ± 1.99 37.77 ± 4.14 53.63 ± 0.39 85.74 ± 0.67 93.51 ± 0.57 76.44 ± 0.62
GraphSAGE 39.93 ± 1.58 3.84$\uparrow$ 43.44 ± 3.19 5.67$\uparrow$ 54.72 ± 0.38 1.09$\uparrow$ 92.19 ± 0.58 6.45$\uparrow$ 96.95 ± 0.41 3.44$\uparrow$ 77.96 ± 0.72 1.52$\uparrow$
GAT 35.62 ± 2.06 39.21 ± 3.08 52.70 ± 0.62 88.75 ± 0.41 93.91 ± 0.35 76.79 ± 0.71
GAT 38.72 ± 1.46 3.10$\uparrow$ 43.44 ± 3.00 4.23$\uparrow$ 54.99 ± 0.71 2.29$\uparrow$ 91.60 ± 0.21 2.85$\uparrow$ 97.76 ± 0.37 3.85$\uparrow$ 79.04 ± 1.27 2.25$\uparrow$
Table 4: Node classification results on large-scale graphs (%). indicates our implementation, while other results are taken from [12, 71]. The top $\mathbf{1^{st}}$, $\mathbf{2^{nd}}$ and $\mathbf{3^{rd}}$ results are highlighted. OOM means out of memory.
ogbn-proteins ogbn-arxiv ogbn-products pokec
# nodes 132,534 169,343 2,449,029 1,632,803
# edges 39,561,252 1,166,243 61,859,140 30,622,564
[…]
GraphGPS 76.83 ± 0.26 70.97 ± 0.41 OOM OOM
GOAT 74.18 ± 0.37 72.41 ± 0.40 82.00 ± 0.43 66.37 ± 0.94
NodeFormer 77.45 ± 1.15 59.90 ± 0.42 72.93 ± 0.13 71.00 ± 1.30
SGFormer 79.53 ± 0.38 72.63 ± 0.13 74.16 ± 0.31 73.76 ± 0.24
NAGphormer 73.61 ± 0.33 70.13 ± 0.55 73.55 ± 0.21 76.59 ± 0.25
Exphormer 74.58 ± 0.26 72.44 ± 0.28 OOM OOM
Polynormer 78.97 ± 0.47 73.46 ± 0.16 83.82 ± 0.11 86.10 ± 0.05
GCN 72.51 ± 0.35 71.74 ± 0.29 75.64 ± 0.21 75.45 ± 0.17
GCN 77.29 ± 0.46 4.78$\uparrow$ 73.60 ± 0.18 1.86$\uparrow$ 82.33 ± 0.19 6.69$\uparrow$ 86.33 ± 0.17 10.88$\uparrow$
GraphSAGE 77.68 ± 0.20 71.49 ± 0.27 78.29 ± 0.16 75.63 ± 0.38
GraphSAGE 82.21 ± 0.32 4.53$\uparrow$ 72.95 ± 0.31 1.46$\uparrow$ 83.89 ± 0.36 5.60$\uparrow$ 85.97 ± 0.21 10.34$\uparrow$
GAT 72.02 ± 0.44 71.95 ± 0.36 79.45 ± 0.59 72.23 ± 0.18
GAT 85.01 ± 0.46 12.99$\uparrow$ 73.30 ± 0.16 1.35$\uparrow$ 80.99 ± 0.16 1.54$\uparrow$ 86.19 ± 0.23 13.96$\uparrow$
Observations on Homophilous Graphs (Table 2). Classic GNNs, with only slight adjustments to hyperparameters, are highly competitive in node classification tasks on homophilous graphs, often outperforming state-of-the-art graph transformers in many cases.
While previously reported results show that most advanced GTs outperform classic GNN on homophilous graphs [12, 71], our implementation of classic GNNs can place within the top two for five datasets, with GCN always ranking in the top three. Specifically, on CS and WikiCS, classic GNNs experience about a 3% accuracy increase, achieving top-three performances. On WikiCS, the accuracy of GAT increases by 3.84%, moving it from seventh to first place, surpassing the leading GT, Polynormer. Similarly, on Computer and Photo, GAT outperforms Polynormer and SGFormer, establishing itself as the top model. On Cora, CiteSeer, PubMed, and Physics, tuning yields significant performance improvements for GCN, with accuracy increases ranging from 1.25% to 3.48%, positioning GCN as the highest-performing model despite its initial lower accuracy compared to advanced GTs.
Observations on Heterophilous Graphs (Table 3). Our implementation has significantly enhanced the previously reported best results of classic GNNs on heterophilous graphs, surpassing specialized GNN models tailored for such graphs and even outperforming the leading graph transformer architectures. This advancement not only supports but also strengthens the findings in [54] that conventional GNNs are strong contenders for heterophilous graphs, challenging the prevailing assumption that they are primarily suited for homophilous graph structures.
The three classic GNNs consistently secure top positions on five out of six heterophilous graphs. Specifically, on well-known page-page networks like Squirrel and Chameleon, our implementation enhances the accuracy of GCN by 4.80% and 5.83% respectively, elevating it to the first place among all models. By comparison, on larger heterophilous graphs such as Amazon-Ratings and Questions, GAT exhibits the highest performance, highlighting the superiority of its message-passing attention mechanism over GTs’ global attention. On Roman-Empire, a 17.68% increase is observed in the performance of GCN. Interestingly, we find that improvements primarily stem from residual connections, which are further analyzed in our ablation study (see Section 5.2).
Observations on Large-scale Graphs (Table 4). Our implementation has significantly enhanced the previously reported results of classic GNNs, with some cases showing double-digit increases in accuracy. It has achieved the best results across these large graph datasets, either homophilous or heterophilous, and has outperformed state-of-the-art graph transformers. This indicates that message passing remains highly effective for learning node representations on large-scale graphs.
Our implementation of classic GNNs demonstrate consistently superior performance, achieving top rankings across all four large-scale datasets included in our study. Notably, GCN emerges as the leading model on ogbn-arxiv and pokec, surpassing all evaluated advanced GTs. Furthermore, on pokec, all three classic GNNs achieve over 10% performance increases by our implementation. For ogbn-proteins, an absolute improvement of 12.99% is observed in the performance of GAT, significantly surpassing SGFormer by 5.48%. Similarly, on ogbn-products, GraphSAGE demonstrates a performance increase, securing the best performance among all evaluated models. In summary, a basic GNN can achieve the best known results on large-scale graphs, suggesting that current GTs have not yet addressed GNN issues such as over-smoothing and long-range dependencies.
5.2 Influence of Hyperparameters on the Performance of GNNs
To examine the unique contributions of different hyperparameters in explaining the enhanced performance of classic GNNs, we conduct a series of ablation analysis by selectively removing elements such as normalization, dropout, residual connections, network depth, and jumping knowledge from GCN, GraphSAGE, and GAT. The effect of these ablations is assessed across homophilous (see Table 5), heterphilous (see Table 6), and large-scale graphs (see Table 7). Our findings, which we detail below, indicate that the ablation of single components affects model accuracy in distinct ways.
Table 5: Ablation study on homophilous graphs (%).
Cora CiteSeer PubMed Computer Photo CS Physics WikiCS
[…]
GCN 85.08 ± 0.52 72.98 ± 0.84 81.32 ± 0.72 93.80 ± 0.29 96.51 ± 0.20 95.80 ± 0.28 97.43 ± 0.05 80.27 ± 0.71
(-) Normalization 85.08 ± 0.52 72.98 ± 0.84 81.32 ± 0.72 93.09 ± 0.13 96.38 ± 0.14 94.47 ± 0.25 97.18 ± 0.15 80.27 ± 0.71
(-) Dropout 82.70 ± 0.87 70.96 ± 1.00 79.50 ± 1.29 93.34 ± 0.11 96.51 ± 0.20 95.59 ± 0.16 97.32 ± 0.06 78.34 ± 0.42
(-) Residual Connections 83.40 ± 1.58 71.60 ± 1.26 80.86 ± 0.58 93.80 ± 0.29 94.68 ± 0.29 94.45 ± 0.12 96.89 ± 0.20 79.38 ± 0.57
(-) Jumping Knowledge 85.08 ± 0.52 72.98 ± 0.84 81.32 ± 0.72 92.97 ± 0.08 95.93 ± 0.32 93.73 ± 0.53 97.00 ± 0.10 79.13 ± 0.91
GraphSAGE 84.18 ± 0.81 71.93 ± 0.85 79.41 ± 0.53 93.59 ± 0.22 96.41 ± 0.17 96.12 ± 0.24 97.21 ± 0.05 80.51 ± 0.48
(-) Normalization 84.18 ± 0.81 71.80 ± 0.19 79.41 ± 0.53 93.07 ± 0.12 96.41 ± 0.17 94.53 ± 0.14 96.73 ± 0.05 80.41 ± 0.70
(-) Dropout 81.54 ± 1.55 70.36 ± 0.47 75.70 ± 1.69 92.71 ± 0.47 96.23 ± 0.10 96.12 ± 0.24 97.14 ± 0.04 80.51 ± 0.57
(-) Residual Connections 80.46 ± 1.19 71.80 ± 0.19 79.41 ± 0.53 93.23 ± 0.26 95.95 ± 0.13 95.87 ± 0.08 97.16 ± 0.04 80.00 ± 0.76
(-) Jumping Knowledge 84.18 ± 0.81 70.00 ± 1.26 79.41 ± 0.53 93.16 ± 0.18 95.82 ± 0.10 94.67 ± 0.26 96.91 ± 0.04 79.89 ± 0.79
GAT 84.64 ± 1.27 72.10 ± 1.10 79.70 ± 0.70 93.93 ± 0.16 96.67 ± 0.13 96.08 ± 0.10 97.30 ± 0.06 80.75 ± 0.74
(-) Normalization 84.64 ± 1.27 72.10 ± 1.10 79.70 ± 0.70 93.86 ± 0.10 96.21 ± 0.17 94.50 ± 0.33 96.99 ± 0.15 80.75 ± 0.74
(-) Dropout 83.18 ± 1.16 70.82 ± 0.76 79.04 ± 0.44 93.35 ± 0.33 96.67 ± 0.13 95.85 ± 0.04 97.30 ± 0.06 80.19 ± 0.16
(-) Residual Connections 83.34 ± 0.24 71.88 ± 0.40 79.70 ± 0.70 93.29 ± 0.31 95.29 ± 0.13 94.47 ± 0.25 96.51 ± 0.05 79.30 ± 0.92
(-) Jumping Knowledge 82.78 ± 0.96 70.36 ± 0.44 79.70 ± 0.70 93.44 ± 0.22 96.43 ± 0.20 94.41 ± 0.28 97.03 ± 0.04 79.56 ± 1.31
Table 6: Ablation study on heterophilous graphs (%).
Squirrel Chameleon Amazon-Ratings Roman-Empire Minesweeper Questions
[…]
GCN 44.50 ± 1.92 46.11 ± 3.16 53.57 ± 0.32 91.35 ± 0.37 97.77 ± 0.38 77.40 ± 1.07
(-) Normalization 43.92 ± 1.59 46.11 ± 3.16 53.57 ± 0.32 91.35 ± 0.37 97.77 ± 0.38 77.40 ± 1.07
(-) Dropout 41.96 ± 1.89 46.11 ± 3.16 50.99 ± 0.21 86.88 ± 0.34 94.33 ± 0.21 76.92 ± 1.55
(-) Residual Connections 44.50 ± 1.92 43.22 ± 3.91 50.31 ± 0.25 75.63 ± 0.48 89.16 ± 0.24 74.49 ± 0.39
(-) Jumping Knowledge 43.10 ± 1.47 43.81 ± 3.96 51.63 ± 0.27 90.73 ± 0.25 95.29 ± 0.62 77.30 ± 0.57
GraphSAGE 39.93 ± 1.58 43.44 ± 3.19 54.72 ± 0.38 92.19 ± 0.58 96.95 ± 0.41 77.96 ± 0.72
(-) Normalization 39.93 ± 1.58 43.44 ± 3.19 54.50 ± 0.38 92.19 ± 0.58 96.95 ± 0.41 76.07 ± 1.56
(-) Dropout 39.93 ± 1.58 40.98 ± 4.65 52.58 ± 0.17 86.50 ± 0.13 93.92 ± 0.73 77.96 ± 0.72
(-) Residual Connections 39.37 ± 2.30 41.24 ± 3.85 53.94 ± 0.40 88.65 ± 0.96 94.41 ± 0.43 76.70 ± 0.87
(-) Jumping Knowledge 38.70 ± 1.95 43.44 ± 3.19 54.27 ± 0.69 90.12 ± 0.54 93.80 ± 0.30 76.80 ± 0.72
GAT 38.72 ± 1.46 43.44 ± 3.00 54.99 ± 0.71 91.60 ± 0.21 97.76 ± 0.37 79.04 ± 1.27
(-) Normalization 38.72 ± 1.46 42.24 ± 3.26 54.99 ± 0.71 90.96 ± 0.20 97.76 ± 0.37 79.02 ± 1.13
(-) Dropout 38.72 ± 1.46 41.08 ± 2.96 53.84 ± 0.87 87.06 ± 0.21 94.10 ± 0.09 78.51 ± 0.36
(-) Residual Connections 38.72 ± 1.46 41.74 ± 2.08 51.24 ± 0.28 79.11 ± 0.88 90.83 ± 0.45 77.07 ± 0.84
(-) Jumping Knowledge 38.72 ± 1.46 41.12 ± 3.94 54.38 ± 0.29 91.60 ± 0.21 93.92 ± 0.73 79.04 ± 1.27
Table 7: Ablation study on large-scale graphs (%).
ogbn-proteins ogbn-arxiv ogbn-products pokec
[…]
GCN 77.29 ± 0.46 73.60 ± 0.18 82.33 ± 0.19 86.33 ± 0.17
(-) Normalization 74.48 ± 1.13 71.97 ± 0.18 80.01 ± 0.48 85.21 ± 0.23
(-) Dropout 74.85 ± 0.87 72.59 ± 0.23 79.30 ± 0.37 84.47 ± 0.38
(-) Residual Connections 73.19 ± 1.46 73.01 ± 0.09 82.33 ± 0.19 79.59 ± 0.97
(-) Jumping Knowledge 77.29 ± 0.46 72.98 ± 0.24 82.33 ± 0.19 86.33 ± 0.17
GraphSAGE 82.21 ± 0.32 72.95 ± 0.31 83.89 ± 0.36 85.97 ± 0.21
(-) Normalization 77.42 ± 0.98 71.52 ± 0.17 82.12 ± 0.31 84.95 ± 0.33
(-) Dropout 80.52 ± 0.49 71.51 ± 0.04 80.36 ± 0.43 83.06 ± 0.28
(-) Residual Connections 81.75 ± 0.53 72.84 ± 0.09 83.89 ± 0.36 85.81 ± 0.45
(-) Jumping Knowledge 82.21 ± 0.32 72.76 ± 0.17 83.89 ± 0.36 85.97 ± 0.21
GAT 85.01 ± 0.46 73.30 ± 0.16 80.99 ± 0.16 86.19 ± 0.23
(-) Normalization 80.32 ± 0.83 71.68 ± 0.36 78.62 ± 0.33 84.63 ± 0.64
(-) Dropout 82.48 ± 0.34 71.83 ± 0.42 77.68 ± 0.21 85.12 ± 0.49
(-) Residual Connections 82.43 ± 0.75 72.76 ± 0.45 80.99 ± 0.16 81.37 ± 0.87
(-) Jumping Knowledge 85.01 ± 0.46 72.94 ± 0.30 80.99 ± 0.16 86.19 ± 0.23
444455556666777788889999101010109393939394949494959595959696969697979797GCN layers (homophilous graphs)Accuracy
444455556666777788889999101010107575757580808080858585859090909095959595GCN layers (heterophilous graphs)ROC-AUC/Accuracy
444455556666777788889999101010107373737373.273.273.273.273.473.473.473.473.673.673.673.673.873.873.873.8GCN layers (large-scale graph)AccuracyComputerPhotoCSRoman-EmpireMinesweeperQuestionsogbn-arxiv
4444555566667777888899991010101093939393949494949595959596969696SAGE layers (homophilous graphs)Accuracy
444455556666777788889999101010107575757580808080858585859090909095959595SAGE layers (heterophilous graphs)ROC-AUC/Accuracy
444455556666777788889999101010107272727272.572.572.572.573737373SAGE layers (large-scale graph)AccuracyComputerPhotoCSRoman-EmpireMinesweeperQuestionsogbn-arxiv
Figure 1: Ablation studies of the number of layers showing, from left to right, results for homophilous graphs, heterophilous graphs, and large-scale graphs, respectively.
Observation 1: Normalization (either BN or LN) is important for node classification on large-scale graphs but less significant on smaller-scale graphs.
Note that we do not conduct an exhaustive search regarding BN and LN, with specific choices detailed in Appendix A. We observe that the ablation of normalization does not lead to substantial deviations on small graphs. However, normalization becomes consistently crucial on large-scale graphs, where its ablation results in accuracy reductions of 3.81% and 4.69% for GraphSAGE and GAT respectively on ogbn-proteins. We believe this is because large graphs display a wider variety of node features, resulting in different data distributions across the graph. Normalization aids in standardizing these features during training, ensuring a more stable distribution.
Observation 2: Dropout is consistently found to be essential for node classification.
Our analysis highlights the crucial role of dropout in maintaining the performance of classic GNNs on both homophilous and heterophilous graphs, with its ablation contributing to notable accuracy declines—for instance, a 3.71% decrease for GraphSAGE on PubMed and a 5.69% decrease on Roman-Empire. This trend persists in large-scale datasets, where the ablation of dropout leads to a 2.44% and 2.53% performance decline for GCN and GAT respectively on ogbn-proteins.
Observation 3: Residual connections can significantly boost performance on specific datasets, exhibiting a more pronounced effect on heterophilous graphs than on homophilous graphs.
While the ablation of residual connections on homophilous graphs does not consistently lead to a significant performance decrease, with observed differences around 2% on Cora, Photo, and CS, the impact is more substantial on large-scale graphs such as ogbn-proteins and pokec. The effect is even more dramatic on heterophilous graphs, with the classic GNNs exhibiting the most significant accuracy reduction on Roman-Empire, for instance, a 15.72% for GCN and 12.49% for GAT. Similarly, on Minesweeper, significant performance drops were observed, emphasizing the critical importance of residual connections, particularly on heterophilous graphs. The complex structures of these graphs often necessitate deeper layers to effectively capture the diverse relationships between nodes. In such contexts, residual connections are essential for model training.
Observation 4: Deeper networks generally lead to greater performance gains on heterophilous graphs compared to homophilous graphs.
As demonstrated in Figure 1, the performance trends for GCN and GraphSAGE are consistent across different graph types. On homophilous graphs and ogbn-arxiv, both models achieve optimal performance with a range of 4 to 8 layers. In contrast, on heterophilous graphs, their performance improves with an increasing number of layers, indicating that deeper networks are more beneficial for these graphs. We discuss scenarios with more than 10 layers in Appendix B.
Observation 5: Jumping knowledge mode has minimal influence on large-scale graphs.
The ablation of this component typically results in marginal accuracy decreases of less than 0.5% for classic GNNs across large-scale graphs, suggesting the lesser impact of jumping knowledge than other components. By comparison, on smaller-scale homophilous graphs, such as CiteSeer and CS, the ablation of jumping knowledge can lead to decreases of around 1.5 to 2% in the model performance. On some heterphilous graphs, the effect is also significant, with the accuracy of GAT decreasing by 2.32% on Chameleon, GraphSAGE by 3.15% and GAT by 3.84% on Minesweeper.
6 Conclusion
Our study provides a thorough reevaluation of the efficacy of foundational GNN models in node classification tasks. Through extensive empirical analysis, we demonstrate that these classic GNN models can reach or surpass the performance of GTs on various graph datasets, challenging the perceived superiority of GTs in node classification tasks. Furthermore, our comprehensive ablation studies provide insights into how various GNN configurations impact performance. We hope our findings promote more rigorous empirical evaluations in graph machine learning research.
Appendix A Datasets and Experimental Details
A.1 Computing Environment
Our implementation is based on PyG [16] and DGL [67]. The experiments are conducted on a single workstation with 8 RTX 3090 GPUs.
A.2 Hyperparameters and Reproducibility
For the hyperparameter selections of classic GNNs, in addition to what we have covered, we list other settings in Tables 8, 9, 10. Notably, for heterophilous graphs, we expand the search range for the number of layers to include three additional settings: $\{11,12,15,20\}$ (See Sec. B.2 for further analysis). This adjustment is based on our empirical evidence suggesting that deep networks tend to yield performance improvements on heterophilous graphs. The ReLU function serves as the non-linear activation. Further details regarding hyperparameters can be found in our code https://github.com/LUOyk1999/tunedGNN.
Due to the large size of the graphs in ogbn-proteins, ogbn-products, and pokec, which prevents full-batch training on GPU memory, we adopt different training strategies. For ogbn-proteins, we utilize the optimized neighbor sampling method [20]. For pokec and ogbn-products, we apply the random partitioning method previously used by GTs [12, 71, 70] to enable mini-batch training. For other datasets, we employ full-batch training. In all experiments, we use the validation set to select the best hyperparameters. GNN denotes our implementation of the GNN model.
Our code is available under the MIT License.
Table 8: Dataset-specific hyperparameter settings of GCN.
Dataset JK ResNet Normalization Dropout rate GNNs layer $L$ Hidden dim LR epoch
Cora False True False 0.5 3 256 0.005 500
Citeseer False True False 0.5 3 256 0.005 500
Pubmed False True False 0.5 2 64 0.01 500
Computer True False LN 0.7 4 512 0.001 1000
Photo True True LN 0.0 9 512 0.001 1000
CS True True LN 0.3 7 512 0.001 1500
Physics True True LN 0.5 9 512 0.001 1500
WikiCS True True False 0.5 10 512 0.001 1000
Squirrel True False BN 0.5 8 64 0.01 500
Chameleon True True BN 0.5 5 256 0.01 200
Amazon-Ratings True True False 0.3 10 512 0.001 2500
Roman-Empire True True False 0.3 11 512 0.001 2500
Minesweeper True True False 0.3 15 512 0.001 2000
Questions True True False 0.2 11 512 0.001 1500
ogbn-proteins False True BN 0.3 3 512 0.01 100
ogbn-arxiv True True BN 0.5 6 256 0.0005 2000
ogbn-products False False LN 0.5 5 256 0.003 300
pokec False True BN 0.2 7 256 0.0005 2000
Table 9: Dataset-specific hyperparameter settings of GraphSAGE.
Dataset JK ResNet Normalization Dropout rate GNNs layer $L$ Hidden dim LR epoch
Cora False True False 0.5 6 256 0.005 500
Citeseer True False False 0.5 5 256 0.001 500
Pubmed False False False 0.5 3 256 0.001 500
Computer True True LN 0.7 7 512 0.001 1000
Photo True True False 0.7 5 512 0.001 1000
CS True True LN 0.0 7 512 0.001 1500
Physics True True LN 0.5 4 512 0.001 1500
WikiCS True True LN 0.5 5 512 0.001 1000
Squirrel True True False 0.0 2 64 0.005 500
Chameleon False True False 0.5 4 256 0.001 200
Amazon-Ratings True True LN 0.3 5 512 0.001 2500
Roman-Empire True True False 0.3 11 512 0.001 2500
Minesweeper True True False 0.3 11 512 0.001 2000
Questions True True LN 0.0 11 512 0.001 1500
ogbn-proteins False True BN 0.3 6 512 0.01 1000
ogbn-arxiv True True BN 0.5 7 256 0.0005 2000
ogbn-products False False LN 0.5 5 256 0.003 1000
pokec False True BN 0.2 7 256 0.0005 2000
Table 10: Dataset-specific hyperparameter settings of GAT.
Dataset JK ResNet Normalization Dropout rate GNNs layer $L$ Hidden dim LR epoch
Cora True True False 0.5 6 256 0.001 500
Citeseer True False False 0.5 4 256 0.001 500
Pubmed False False False 0.5 2 256 0.001 500
Computer True True LN 0.7 6 512 0.001 1000
Photo True True LN 0.0 5 512 0.001 1000
CS True True LN 0.3 5 512 0.001 1500
Physics True True LN 0.0 5 512 0.001 1500
WikiCS True True False 0.5 7 512 0.001 1000
Squirrel False False False 0.0 3 256 0.001 500
Chameleon True True BN 0.5 4 256 0.005 200
Amazon-Ratings True True False 0.3 7 512 0.001 2500
Roman-Empire False True LN 0.3 6 512 0.001 2500
Minesweeper True True False 0.3 15 512 0.001 2000
Questions False True LN 0.2 4 512 0.001 1500
ogbn-proteins False True BN 0.3 7 512 0.01 1000
ogbn-arxiv True True BN 0.5 6 256 0.0005 2000
ogbn-products False False LN 0.5 5 256 0.003 1000
pokec False True BN 0.2 7 256 0.0005 2000
Table 11: Node classification results on ogbn-proteins (%).
ogbn-proteins
[…]
DeeperGCN 85.80 ± 0.17
GAT (with edge features) 87.82 ± 0.16
Appendix B Additional Benchmark Results
B.1 GAT with Edge Features on ogbn-proteins
While DeepGCN [33] introduced training models up to 56 layers deep and DeeperGCN [34] further extended this to 112 layers, our experiments suggest that such depth is not necessary. Specifically, while the DeeperGCN achieved an accuracy of 85.50% on ogbn-proteins, it utilized edge features as input, a configuration not commonly employed in the standard baselines of the OGB dataset [24]. As our experiments do not incorporate edge features on ogbn-proteins, we exclude DeeperGCN from our main text to maintain a fair comparison.
Now we incorporate edge features into the GAT, same as the approach in [68], with the results shown in Table 11. A 6-layer GAT achieve an accuracy of 87.47%, significantly surpassing the 85.50% by DeeperGCN. This demonstrates that GNNs do not need to be as deep as proposed by DeeperGCN; a range of 2 to 10 layers is typically sufficient.
B.2 Deeper Networks on Heterophilous Graphs
On heterophilous graphs, the performance of classic GNNs improves with an increasing number of layers limited to 10, as evidenced by Figure 1 in the main text. We explore scenarios with more than 10 layers in this subsection. Specifically, we consider GCN and GAT with layer configurations of 11, 12, 15, and 20 for the Amazon-Ratings, Roman-Empire, and Minesweeper datasets. The results are shown in Table 12. The variation in the optimal number of layers ($L$) could stem from the distinct structures inherent in different graphs. Heterophilous graphs may have more complex structures, thus necessitating a higher $L$. However, the slight improvements observed with larger $L$ values suggest that very deep networks may not yield significantly better results. Overall, the best results for classic GNNs are achieved when $L$ is limited to 15.
Table 12: Ablation study of the number of layers $L$ on heterophilous graphs (%).
Amazon-Ratings Roman-Empire Minesweeper
[…]
GCN ($L=11$) 53.52 ± 0.19 91.35 ± 0.37 97.01 ± 0.18
GCN ($L=12$) 53.66 ± 0.11 90.75 ± 0.37 97.16 ± 0.28
GCN ($L=15$) 53.65 ± 1.02 91.10 ± 0.41 97.77 ± 0.38
GCN ($L=20$) 53.50 ± 0.58 91.10 ± 0.76 97.76 ± 0.30
GAT ($L=11$) 54.61 ± 0.21 91.07 ± 0.40 97.11 ± 0.09
GAT ($L=12$) 54.55 ± 0.18 90.79 ± 0.44 97.59 ± 0.15
GAT ($L=15$) 55.01 ± 0.45 90.98 ± 0.41 97.76 ± 0.37
GAT ($L=20$) 55.01 ± 0.17 91.30 ± 0.38 97.56 ± 0.47
Appendix C Limitations & Broader Impacts
Broader Impacts. This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Limitations. In this study, we focus solely on the node classification task, without delving into graph classification and link prediction tasks. It would be beneficial to extend our benchmarking efforts to include classic GNNs in graph-level and edge-level tasks. Additionally, we did not conduct an exhaustive hyperparameter search due to the large set of hyperparameters and datasets. Therefore, future research could potentially uncover more favourable results by undertaking a more comprehensive exploration of the hyperparameter space.